OpenStudy (agl202):

Quick help... Consider the function f(x) = x2 − 2x + 6. Find the value of f(x) when x = 3 and select the correct answer below. 1 3 6 9

OpenStudy (anonymous):

Plug in x

OpenStudy (anonymous):

All you need to do is plug 3 in wherever you see an x and then solve. Does that clear it up?

OpenStudy (jhannybean):

\(f(3) = (3)^2-2(3)+6\)

OpenStudy (agl202):

it's 9, right?

OpenStudy (solomonzelman):

For example, if I needed to find the value of f(x) when x=4, THEN this is what I would do: \(\large\color{black}{ \displaystyle f(x)=x^2-2x+6 }\) \(\large\color{black}{ \displaystyle f(\color{red}{4})=\color{red}{(4)}^2-2\color{red}{(4})+6 }\) \(\large\color{black}{ \displaystyle f(\color{red}{4})=16-8+6 }\) \(\large\color{black}{ \displaystyle f(\color{red}{4})=14 }\) ----------------------------------------------- Yes, 9 is right for x=3

OpenStudy (anonymous):

Yep, it is 9 =)

OpenStudy (agl202):

Duh! Lol, Thx guys! :D

OpenStudy (solomonzelman):

Anytime....